# PhysDepth: Plug-and-Play Physical Refinement for Monocular Depth Estimation in Challenging Environments

### v2

@@ 1 Introduction
Monocular depth estimation (MDE) is a fundamental task for autonomous driving [park2021pseudo, wang2019pseudo] and robotics [zhou2019does, dong2022towards]. While recent self-supervised methods have advanced significantly, their performance remains inherently limited by a primarily data-driven paradigm [gasperini2023robust, rajapaksha2024deep, liu2021self, vankadari2020unsupervised, wang2021regularizing, vankadari2023sun, zheng2023steps, godard2019digging, vankadari2024dusk, spencer2020defeat]. This reliance forces models to learn brittle, non-physical heuristics (e.g., ’finer texture: closer’ or ’brighter lights: closer’) rather than robust physical principles. These spurious correlations often struggle in challenging environments, resulting in suboptimal results.
@@
Figure 1: Plug-and-Play of the PhysDepth. we show that adding PhysDepth to a baselines like MonoDepth2 [godard2019digging], md4ALL[gasperini2023robust], and RNW [wang2021regularizing] effectively corrects common failure modes.
@@
As illustrated in Figure 1, SOTA MDE models, including MonoDepth2 [godard2019digging], md4ALL[gasperini2023robust], and RNW [wang2021regularizing] fail in challenging environments, producing hazardous artifacts from real-world conditions like road reflections and glare.
@@ 2 Related Work
Depth estimation under challenging conditions (e.g., rain or night) has attracted increased attention recently. Some SOTA methods focus on feature representation and regularizers. For example, Zhang et al. [zhang2025lightweight] leverages CoMoGAN to transform daytime images into challenging scenes, thereby circumventing the need for a separate nighttime dataset. Yan et al. [Yan_2025_CVPR] also use synthetic adverse data to train their model. Zhang et al [Zhang_2025_ICCV] use CLIP and DINO under contrastive language guidance for dark images. Chung et al. [Chung_2025_ICCV] proposed a test-time adaptation for challenging condition depth estimation. DeFeat-Net [spencer2020defeat] uses a cross-domain dense feature representation. RNW [wang2021regularizing] uses image enhancement and a GAN-based regularizer. Cong et al. [10696933] introduce a structure-regularized framework cropped multi-scale consistency loss for low visibility and illumination conditions. Meanwhile, other SOTA methods focus on domain adaptation networks. ADIDS [liu2021self] uses two networks for day-time and nighttime images. Vankadari et al. [vankadari2020unsupervised] and ITDFA [zhao2022unsupervised] both adapted a domain adaptation network trained on daytime images to function in dark environments. Gasperini et al. [gasperini2023robust] also convert challenging environment images to daytime conditions using a translation model. WSGD [vankadari2023sun] adds denoising into domain adaptation network. Hou et al. [10539960] introduce an illumination compensation PoseNet for lighting adaptation, and a cross-layer adaptive fusion module to enhance feature complementarity. Sharma et al. [sharma2020nighttime] developed a translation network to create nighttime stereo images from daytime versions, then trained on a stereo network. Also, some SOTA methods use additional information. Lu et al. [lu2021alternative] combined thermal and RGB data to enhance the accuracy of nighttime depth estimation. R4dyn [gasperini2021r4dyn] uses radar at nighttime.
@@ 3.1 Empirical Study, Motivation, and Base Model
We argue that existing SOTA methods overlook robust physical information. To prove this, we conduct an empirical study that measures the covariance $Cov$ between a model’s error metric and a controlled, physically-plausible injection of atmosphere attenuation $\beta$ [he2010single].
This study uses a test set containing challenging condition images $I$ and corresponding ground-truth depth maps $D_{gt}$. We generate a sequence of perturbed images by applying the volumetric scattering model [he2010single, tan2008visibility, 10.1145/1399504.1360671, narasimhan2000chromatic, narasimhan2002vision] in Equation 1 :
@@
Interpretation of the covariance. As shown in Figure 2, Our PhysDepth (purple solid line) achieves the lowest AbsRel error across the entire range of attenuation coefficients, consistently and significantly outperforming all competing models. As $\beta$ increases, the error of all models rises. However, PhysDepth exhibits the flattest performance curve, indicating that its performance degrades the least as conditions worsen. This superior stability is also quantified by the $Cov$ score shown in the legend; our method achieves the best (lowest) score of -0.040, while all other methods have positive (worse) scores, confirming its exceptional robustness to challenging environments. Also, PhysDepth’s $Cov<0$ means that it has successfully learned the physical priors, treating attenuation as a valid, independent signal for depth. SOTAs such as STEPS [zheng2023steps], DepthDark [zeng2025depthdark], md4ALL-AD [gasperini2023robust], md4ALL-DD [gasperini2023robust], their $Cov\approx 0.0$ (No Correlation) means each model’s performance is decoupled from attenuation, indicating it only rely on other cues, such as semantic context or geometry. For other methods like WSGB [vankadari2023sun] and RNW [wang2021regularizing], their $Cov>0$ (Positive Correlation) indicates that each model treats attenuation as noise that occludes its primary cues. This empirical study motivates us to design PhysDepth based on robust physical information.
Figure 3 illustrates the design of PhysDepth. We select a modern Transformer-based architecture as the encoder in our primary Base MDE Models (Figure 3 - b). We also follow Piccinelli et al. [piccinelli2024unidepth] to employ both ConvNeXt-T [liu2022convnet] and MiDaS [ranftl2020towards] to serve as the encoder in the base model.
@@ 3.2 Physical Prior Module
We start by conducting experiments to verify the relationships between image channels and the accuracy of MDE results. Particularly, RGB images from RobotCar-Night dataset [RobotCarDatasetIJRR] are replaced by their single-channel images and two-channel images for comparison, respectively. We use the base MDE model, Pose Network, and Photometric Loss in Figure 3 as a baseline. Table 1 shows that single-channel image settings outperform their two-channel counterparts, surpassing RGB image settings. Notably, the red channel setting (Baseline+R) yields the best results.
Table 1: Using Different Channel Images from RobotCar-Night Dataset [RobotCarDatasetIJRR]: R means Red Channel, so do B (Blue Channel) and G (Green Channel), RGB. The best results are in bold. Base MDE model uses ViT-B/.
@@
The red channel’s superior performance attributes to a two-fold physical advantage: 1) Rayleigh Scattering: light is scattered by the atmospheric medium itself, even in clear air. This law states that the shorter-wavelength blue and green channels are scattered more than the red channel, degrading their signal integrity over distance. 2) Signal-to-Noise Ratio (SNR): This fundamental scattering advantage is compounded by the realities of challenging-condition illumination. Artificial lighting (e.g., brake lights, streetlights) is mainly red/yellow. As a result, the sensor’s red channel receives a stronger and cleaner signal, while the blue and green channels are more likely to capture sensor noise. Furthermore, combining multiple channels can introduce additional noise [hong2024you], leading to inaccurate MDE results. Thus, we use red channel images for PhysDepth.
@@
To implement PPM, we employ a modern hierarchical backbone, ConvNeXt [liu2022convnet], to extract image features and establish the correlation between red channel values and depth values (Figure 3 - a). To enhance the robustness of the Base MDE Model, we inject the physical features from the PPM at multiple scales. The intermediate feature maps from the ConvNeXt encoder stages (Stage 2, Stage 3, Stage 4) are fused with the feature maps of the Base MDE Decoder at corresponding resolutions via channel-wise concatenation.
@@ 3.3 Red Channel Attenuation Loss (RCA loss)
We then use gradient Spherical Gaussian (SG) [li2022physically, wang2009all, wang2021learning] to model $I_{0}$ in Equation 5, where $\lambda\in\mathbb{R}$ is a learnable parameter. It is because the naive application of the Beer-Lambert law would require solving for the per-pixel intrinsic brightness $I_{0}$, which is ill-posed from a single image. We instead simplify this term by modeling its contribution as a single, globally learnable parameter $\lambda$. We are therefore not claiming to model the true $I_{0}$ of every pixel. Rather, the term $\frac{1}{\mu}(g\lambda-1)$ functions as a learned scalar bias in the regression. This simplification makes the problem tractable and prevents the model from learning a trivial identity by encoding depth in the $I_{0}$ term.
@@
RCA loss serves as a self-supervised signal to guide the training processes of PPM. Besides RCA loss, we also train our model with another supervisory signal: photometric reconstruction loss function $\mathcal{L}_{\text{p}}$ from SfMLearner [zhou2017unsupervised]. It requires four inputs as Figure 3 shows: Image $R_{t},R_{t-1}$, Transformation $T_{t\rightarrow t-1}$, predicted depth map $D_{t}$, and camera intrinsics matrix $K$. We add the weak velocity supervision $\mathcal{L}_{\text{v}}$ [guizilini20203d] to address scale ambiguity and ensure consistent predictions of photometric reconstruction [gasperini2023robust]. Also following [godard2019digging, gasperini2023robust], we consider partial occlusions via the minimum reprojection error $\mathcal{L}_{\text{p}}$, and we ignore static pixels. Another loss $\mathcal{L}_{\text{s}}$ is used to promote smoothness and preserves edges [godard2017unsupervised] The total loss function is $\mathcal{L}=\mathcal{L}_{\text{p}}+\mathcal{L}_{2}+\mathcal{L}_{\text{v}}+\mathcal{L}_{\text{s}}$.
@@
Table 2: Quantitative Results - RobotCar-Night [RobotCarDatasetIJRR]: We scale depth up to 50 meters as same as [gasperini2023robust, vankadari2023sun]. The best results are in bold for each metric. The results presented here are without post-processing. AD and $D$ mean different training strategies for md4All.
@@
MonoDepth2 [godard2019digging] 0.453 11.420 0.700
DeFeat-Net [spencer2020defeat] 0.334 8.606 0.586
ADDS [liu2021self] 0.287 7.985 0.490
DepthSegNet24 [thanh2024depthsegnet24] 0.195 5.936 0.734
ADFA [vankadari2020unsupervised] 0.231 8.800 0.620
RNW [wang2021regularizing] 0.185 6.549 0.733
WSGD [vankadari2023sun] 0.174 6.302 0.754
STEPS [zheng2023steps] 0.170 6.797 0.758
DepthDark [zeng2025depthdark] 0.157 4.284 0.760
md4All-AD [gasperini2023robust] 0.124 3.612 0.858
md4All-DD [gasperini2023robust] 0.122 3.604 0.849
Syn2Real-Depth [yan2025synthetic] 0.110 3.455 0.862
@@
Table 3: Quantitative Results - nuScenes-Night [caesar2020nuscenes]: We scale depth up to 80 meters as same as [gasperini2021r4dyn, gasperini2023robust]. The best results are in bold for each metric. All results presented here are without post-processing. AD and $D$ mean different training strategies for md4All.
@@
STEPS [zheng2023steps] 0.474 13.158 0.334
SRNSD [cong2024srnsd] 0.352 11.942 0.419
Robust-Depth [saunders2023self] 0.276 10.470 0.607
RNW [wang2021regularizing] 0.290 11.289 0.557
R4Dyn [gasperini2021r4dyn] 0.219 10.542 0.623
DepthDark [zeng2025depthdark] 0.210 7.764 0.630
Syn2Real-Depth [yan2025synthetic] 0.171 7.689 0.732
MonoDepth2 [godard2019digging] 0.147 6.941 0.827
md4All-AD [gasperini2023robust] 0.182 6.372 0.753
md4All-DD [gasperini2023robust] 0.128 6.607 0.845
@@
Table 4: Quantitative Results - nuScenes-Rain [caesar2020nuscenes]: We scale depth up to 80 meters as same as [gasperini2021r4dyn, gasperini2023robust]. The best results are in bold for each metric. All results presented here are without post-processing. AD and $D$ mean different training strategies for md4All.
@@
RNW [vankadari2020unsupervised] 0.295 9.341 0.572
R4Dyn [gasperini2021r4dyn] 0.243 10.055 0.570
AdaBins [bhat2021adabins] 0.234 7.088 0.616
md4All-DD [gasperini2023robust] 0.214 8.376 0.680
MonoDepth [liu2021self] 0.157 7.453 0.795
md4All-AD [gasperini2023robust] 0.141 6.915 0.830
Syn2Real-Depth [yan2025synthetic] 0.133 6.926 0.817
@@ 4 Experiments
Implementation Details: Our model is built using PyTorch [paszke2019pytorch]. We train our model with a batch size of 4, Adam optimizer [kingma2014adam] is used with $\beta_{1}=0.9$ and $\beta_{2}=0.999$, we set the learning rate to $0.0002$. Meanwhile, we follow Zhou et al. [zhou2017unsupervised], which we multiply the predicted depth maps by a scalar $\hat{s}$ to match the median with the groundtruth, i.e., $\hat{s}=\operatorname{median}(D_{\text{gt}})/\operatorname{median}(D_{\text{pred}})$.
Datasets: We evaluate PhysDepth on three datasets (RobotCar-Night [RobotCarDatasetIJRR], nuScenes-Night [caesar2020nuscenes], and nuScenes-Rain [caesar2020nuscenes])
@@
We evaluate our proposed model both quantitatively and qualitatively. For quantitative analysis, we perform the comparison against SOTA methods on all three datasets mentioned above. We use three metrics, same as those used in [eigen2014depth, godard2019digging] - Abs Rel, RMSE, and $\delta<1.25$. For qualitative analysis, we compare estimated depth maps from PhysDepth and SOTA models using images from RobotCar-Night and nuScenes (night and rain).

### v3

@@ 1 Introduction
Monocular depth estimation (MDE) is a fundamental task for autonomous driving [31, 51] and robotics [61, 7]. While recent self-supervised methods have advanced significantly, their performance remains inherently limited by a primarily data-driven paradigm [12, 36, 24, 45, 49, 46, 60, 15, 47, 41]. This reliance forces models to learn brittle, non-physical heuristics (e.g., ’finer texture: closer’ or ’brighter lights: closer’) rather than robust physical principles. These spurious correlations often struggle in challenging environments, resulting in suboptimal results.
@@
Figure 1: Plug-and-Play of the PhysDepth. we show that adding PhysDepth to a baselines like MonoDepth2 [15], md4ALL[12], and RNW [49] effectively corrects common failure modes.
@@
As illustrated in Figure 1, SOTA MDE models, including MonoDepth2 [15], md4ALL[12], and RNW [49] fail in challenging environments, producing hazardous artifacts from real-world conditions like road reflections and glare.
@@ 2 Related Work
Depth estimation under challenging conditions (e.g., rain or night) has attracted increased attention recently. Some SOTA methods focus on feature representation and regularizers. For example, Zhang et al. [57] leverages CoMoGAN to transform daytime images into challenging scenes, thereby circumventing the need for a separate nighttime dataset. Yan et al. [54] also use synthetic adverse data to train their model. Zhang et al [58] use CLIP and DINO under contrastive language guidance for dark images. Chung et al. [4] proposed a test-time adaptation for challenging condition depth estimation. DeFeat-Net [41] uses a cross-domain dense feature representation. RNW [49] uses image enhancement and a GAN-based regularizer. Cong et al. [5] introduce a structure-regularized framework cropped multi-scale consistency loss for low visibility and illumination conditions. Meanwhile, other SOTA methods focus on domain adaptation networks. ADIDS [24] uses two networks for day-time and nighttime images. Vankadari et al. [45] and ITDFA [59] both adapted a domain adaptation network trained on daytime images to function in dark environments. Gasperini et al. [12] also convert challenging environment images to daytime conditions using a translation model. WSGD [46] adds denoising into domain adaptation network. Hou et al. [19] introduce an illumination compensation PoseNet for lighting adaptation, and a cross-layer adaptive fusion module to enhance feature complementarity. Sharma et al. [39] developed a translation network to create nighttime stereo images from daytime versions, then trained on a stereo network. Also, some SOTA methods use additional information. Lu et al. [26] combined thermal and RGB data to enhance the accuracy of nighttime depth estimation. R4dyn [11] uses radar at nighttime.
@@ 3.1 Empirical Study, Motivation, and Base Model
We argue that existing SOTA methods overlook robust physical information. To prove this, we conduct an empirical study that measures the covariance $Cov$ between a model’s error metric and a controlled, physically-plausible injection of atmosphere attenuation $\beta$ [17].
This study uses a test set containing challenging condition images $I$ and corresponding ground-truth depth maps $D_{gt}$. We generate a sequence of perturbed images by applying the volumetric scattering model [17, 42, 9, 29, 30] in Equation 1 :
@@
Interpretation of the covariance. As shown in Figure 2, Our PhysDepth (purple solid line) achieves the lowest AbsRel error across the entire range of attenuation coefficients, consistently and significantly outperforming all competing models. As $\beta$ increases, the error of all models rises. However, PhysDepth exhibits the flattest performance curve, indicating that its performance degrades the least as conditions worsen. This superior stability is also quantified by the $Cov$ score shown in the legend; our method achieves the best (lowest) score of -0.040, while all other methods have positive (worse) scores, confirming its exceptional robustness to challenging environments. Also, PhysDepth’s $Cov<0$ means that it has successfully learned the physical priors, treating attenuation as a valid, independent signal for depth. SOTAs such as STEPS [60], DepthDark [56], md4ALL-AD [12], md4ALL-DD [12], their $Cov\approx 0.0$ (No Correlation) means each model’s performance is decoupled from attenuation, indicating it only rely on other cues, such as semantic context or geometry. For other methods like WSGB [46] and RNW [49], their $Cov>0$ (Positive Correlation) indicates that each model treats attenuation as noise that occludes its primary cues. This empirical study motivates us to design PhysDepth based on robust physical information.
Figure 3 illustrates the design of PhysDepth. We select a modern Transformer-based architecture as the encoder in our primary Base MDE Models (Figure 3 - b). We also follow Piccinelli et al. [35] to employ both ConvNeXt-T [25] and MiDaS [37] to serve as the encoder in the base model.
@@ 3.2 Physical Prior Module
We start by conducting experiments to verify the relationships between image channels and the accuracy of MDE results. Particularly, RGB images from RobotCar-Night dataset [27] are replaced by their single-channel images and two-channel images for comparison, respectively. We use the base MDE model, Pose Network, and Photometric Loss in Figure 3 as a baseline. Table 1 shows that single-channel image settings outperform their two-channel counterparts, surpassing RGB image settings. Notably, the red channel setting (Baseline+R) yields the best results.
Table 1: Using Different Channel Images from RobotCar-Night Dataset [27]: R means Red Channel, so do B (Blue Channel) and G (Green Channel), RGB. The best results are in bold. Base MDE model uses ViT-B/.
@@
The red channel’s superior performance attributes to a two-fold physical advantage: 1) Rayleigh Scattering: light is scattered by the atmospheric medium itself, even in clear air. This law states that the shorter-wavelength blue and green channels are scattered more than the red channel, degrading their signal integrity over distance. 2) Signal-to-Noise Ratio (SNR): This fundamental scattering advantage is compounded by the realities of challenging-condition illumination. Artificial lighting (e.g., brake lights, streetlights) is mainly red/yellow. As a result, the sensor’s red channel receives a stronger and cleaner signal, while the blue and green channels are more likely to capture sensor noise. Furthermore, combining multiple channels can introduce additional noise [18], leading to inaccurate MDE results. Thus, we use red channel images for PhysDepth.
@@
To implement PPM, we employ a modern hierarchical backbone, ConvNeXt [25], to extract image features and establish the correlation between red channel values and depth values (Figure 3 - a). To enhance the robustness of the Base MDE Model, we inject the physical features from the PPM at multiple scales. The intermediate feature maps from the ConvNeXt encoder stages (Stage 2, Stage 3, Stage 4) are fused with the feature maps of the Base MDE Decoder at corresponding resolutions via channel-wise concatenation.
@@ 3.3 Red Channel Attenuation Loss (RCA loss)
We then use gradient Spherical Gaussian (SG) [22, 48, 52] to model $I_{0}$ in Equation 5, where $\lambda\in\mathbb{R}$ is a learnable parameter. It is because the naive application of the Beer-Lambert law would require solving for the per-pixel intrinsic brightness $I_{0}$, which is ill-posed from a single image. We instead simplify this term by modeling its contribution as a single, globally learnable parameter $\lambda$. We are therefore not claiming to model the true $I_{0}$ of every pixel. Rather, the term $\frac{1}{\mu}(g\lambda-1)$ functions as a learned scalar bias in the regression. This simplification makes the problem tractable and prevents the model from learning a trivial identity by encoding depth in the $I_{0}$ term.
@@
RCA loss serves as a self-supervised signal to guide the training processes of PPM. Besides RCA loss, we also train our model with another supervisory signal: photometric reconstruction loss function $\mathcal{L}_{\text{p}}$ from SfMLearner [62]. It requires four inputs as Figure 3 shows: Image $R_{t},R_{t-1}$, Transformation $T_{t\rightarrow t-1}$, predicted depth map $D_{t}$, and camera intrinsics matrix $K$. We add the weak velocity supervision $\mathcal{L}_{\text{v}}$ [16] to address scale ambiguity and ensure consistent predictions of photometric reconstruction [12]. Also following [15, 12], we consider partial occlusions via the minimum reprojection error $\mathcal{L}_{\text{p}}$, and we ignore static pixels. Another loss $\mathcal{L}_{\text{s}}$ is used to promote smoothness and preserves edges [14] The total loss function is $\mathcal{L}=\mathcal{L}_{\text{p}}+\mathcal{L}_{2}+\mathcal{L}_{\text{v}}+\mathcal{L}_{\text{s}}$.
@@
Table 2: Quantitative Results - RobotCar-Night [27]: We scale depth up to 50 meters as same as [12, 46]. The best results are in bold for each metric. The results presented here are without post-processing. AD and $D$ mean different training strategies for md4All.
@@
MonoDepth2 [15] 0.453 11.420 0.700
DeFeat-Net [41] 0.334 8.606 0.586
ADDS [24] 0.287 7.985 0.490
DepthSegNet24 [43] 0.195 5.936 0.734
ADFA [45] 0.231 8.800 0.620
RNW [49] 0.185 6.549 0.733
WSGD [46] 0.174 6.302 0.754
STEPS [60] 0.170 6.797 0.758
DepthDark [56] 0.157 4.284 0.760
md4All-AD [12] 0.124 3.612 0.858
md4All-DD [12] 0.122 3.604 0.849
Syn2Real-Depth [55] 0.110 3.455 0.862
@@
Table 3: Quantitative Results - nuScenes-Night [3]: We scale depth up to 80 meters as same as [11, 12]. The best results are in bold for each metric. All results presented here are without post-processing. AD and $D$ mean different training strategies for md4All.
@@
STEPS [60] 0.474 13.158 0.334
SRNSD [6] 0.352 11.942 0.419
Robust-Depth [38] 0.276 10.470 0.607
RNW [49] 0.290 11.289 0.557
R4Dyn [11] 0.219 10.542 0.623
DepthDark [56] 0.210 7.764 0.630
Syn2Real-Depth [55] 0.171 7.689 0.732
MonoDepth2 [15] 0.147 6.941 0.827
md4All-AD [12] 0.182 6.372 0.753
md4All-DD [12] 0.128 6.607 0.845
@@
Table 4: Quantitative Results - nuScenes-Rain [3]: We scale depth up to 80 meters as same as [11, 12]. The best results are in bold for each metric. All results presented here are without post-processing. AD and $D$ mean different training strategies for md4All.
@@
RNW [45] 0.295 9.341 0.572
R4Dyn [11] 0.243 10.055 0.570
AdaBins [2] 0.234 7.088 0.616
md4All-DD [12] 0.214 8.376 0.680
MonoDepth [24] 0.157 7.453 0.795
md4All-AD [12] 0.141 6.915 0.830
Syn2Real-Depth [55] 0.133 6.926 0.817
@@ 4 Experiments
Implementation Details: Our model is built using PyTorch [33]. We train our model with a batch size of 4, Adam optimizer [21] is used with $\beta_{1}=0.9$ and $\beta_{2}=0.999$, we set the learning rate to $0.0002$. Meanwhile, we follow Zhou et al. [62], which we multiply the predicted depth maps by a scalar $\hat{s}$ to match the median with the groundtruth, i.e., $\hat{s}=\operatorname{median}(D_{\text{gt}})/\operatorname{median}(D_{\text{pred}})$.
Datasets: We evaluate PhysDepth on three datasets (RobotCar-Night [27], nuScenes-Night [3], and nuScenes-Rain [3])
@@
We evaluate our proposed model both quantitatively and qualitatively. For quantitative analysis, we perform the comparison against SOTA methods on all three datasets mentioned above. We use three metrics, same as those used in [8, 15] - Abs Rel, RMSE, and $\delta<1.25$. For qualitative analysis, we compare estimated depth maps from PhysDepth and SOTA models using images from RobotCar-Night and nuScenes (night and rain).